1. What is the electric field 8 m from the center of the terminal ofa Van de Graaff with a 2.13 mC charge, noting that the field is equivalent to that of a point charge at the center of the terminal? Give answer in terms of 106 N/C.

2. How strong is the attractive force between a glass rod with a 0.616 μC charge and a silk cloth with a –0.828 μC charge, which are 9.9 cm apart, using the approximation that they act like point charges? Hint: your answer should be positive.

3. Common static electricity involves charges ranging from nanocoulombs to microcoulombs. How many electrons are needed to form a charge of -4.27 nC? Give your answer in terms of 1010 electrons.

4. What magnitude point charge creates a 5,156 N/C electric field at a distance of 0.86 m? Give answer in terms of 10-8.

5. Common static electricity involves charges ranging from nanocoulombs to microcoulombs. How many electrons must be removed from a neutral object to leave a net charge difference of 0.83 μC? Give your answer in terms of 1012 electrons.
Answer #1

1) We know that electric field due to a point charge q at a distance r is given by

E=kq/r^2 where k=9×10^(9) Nm2/C2

Here, r=8m, q=2.13mC

E= (9×10^9)×(2.13×10^-3)/(8^2)

E= 2.995×10^6 N/C